I am 17, My dad is 49, how old does he need to be twice my age.
***I just need the equation***

Answer by jim_thompson5910(35256)

Let x = number of years need until he is twice your age

So if you add x years to your 17 years of age, you will be 17+x years old (in x years from now)

He will be 49+x years old (x years from now)

You want him to be twice your age, so this means

(His age x years from now) = 2*(your age x years from now)

49+x = 2(17+x)

and that's your initial equation. I won't solve it for you because you just wanted the equation, but I figured the answer doesn't hurt (so you can check). The solution to the equation above is x = 15. This means that in 15 years, he will be twice your age. So in 15 years you will be 17+15 = 32 and he will be 49+15 = 64.